Calculate the molar mass of barium citrate.


1
Expert's answer
2021-04-09T04:05:20-0400

Ba(NO3)2

Molar mass of Ba = 137.327

Molar Mass of N = 14.0067

= 14.0067 × 2 = 28.0134

Molar Mass of O = 15.999

= 15.999 × 6 = 95.994

= 137.327 + 28.0134 + 95.994

= 261.3344g/mol